OpenStudy (anonymous):

need help! will give medal! :D <3

OpenStudy (anonymous):

With ?

OpenStudy (anonymous):

hold on trying to upload it

OpenStudy (anonymous):

i gtg now :(

OpenStudy (anonymous):

sowiie v_v

OpenStudy (anonymous):

that's ok its not uploading any ways lol thank u though! :D

OpenStudy (anonymous):

if you can do this for me I would greatly apprecaiate it! thank you! :D

OpenStudy (anonymous):

ill try...

OpenStudy (anonymous):

ok what do need me to do?

OpenStudy (anonymous):

well what the attachments says to do

OpenStudy (anonymous):

oh yeah...........duh

OpenStudy (anonymous):

whew...... that's gonna take me a while,hold on.........

OpenStudy (anonymous):

lol!

OpenStudy (anonymous):

kk lol thanks though for doing it for me!

OpenStudy (anonymous):

look at the attachments @leewilson

OpenStudy (anonymous):

@johnweldon1993 can u do it?

OpenStudy (johnweldon1993):

So basically you just need to go around your house and find some rectangular prisms (think of cereal boxes...iphone boxes (anything like that but rectangular lol) and then some circles...(a bowl...a paper plate...normal plate etc...) Then take like a tape measure....and measure the length width and height for the prisms....and the radius (half the diameter) of the circles... then just record and then solve those equations...

OpenStudy (anonymous):

thanks @johnweldon1993

OpenStudy (johnweldon1993):

So \[\large \text{Volume of Rectangular Prism} = Length \times Width \times Height\] \[\large \text{Surface Area of Rectangular Prism} = 2(WH \times LW \times LH)\] \[\large \text{Circumference of circle} = 2\pi r\] \[\large \text{Area of circle} = \pi r^2\]

OpenStudy (anonymous):

is there a volume off a circle? @johnweldon1993

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

volume of a circle ? lmao

OpenStudy (anonymous):

lol nooo I need help with the attachment can you do it and than give it to me?

OpenStudy (anonymous):

I really don't get what they are asking you to do. Are they asking you to show everything about each of the cubic shapes (or volume containing shapes, such as Sphere and other prisms like rectangular prism, triangular prism and all that) ?

OpenStudy (anonymous):

You can ask a question about each thing in a chart, and I'll explain it to you, but I am not here to do your hw.

OpenStudy (anonymous):

I can do it, but I am not willing to just give it to you without having you learn something. If anyone else is willing to do that, then good luck. If you just want \( \bf HELP\) then I am willing to \( \bf HELP\).

OpenStudy (anonymous):

please I am behind I am trying to catch up. just this one time! I will greatly appreciate it. it would help me so much @GABI_J

OpenStudy (anonymous):

I don't know what you are exactly doing. Ask em questions like "how to find the area of a ____ " and I'll explain. we will do each cell in the chart like this, it won't take long.

OpenStudy (anonymous):

what is the dimensions of a paper plate that has a 7in length 2in width and 2in height

OpenStudy (anonymous):

@GABI_J